1. Jose spends her income $40 on two goods. Good X and Good Y. The price of good X is $8 and price of good Y is $2.



(a) Draw the budget constraint to show how Jose can spend his money on the two goods.



(b) Calculate the slope of the budget line.



(c) If the price of good X increase by $2 what will happen to the budget line?



(d) If the price of good Y decreases by $1 what will happen to the budget line?

1
Expert's answer
2022-01-20T09:53:30-0500

Solution:

a.). Budget constraint: M = PxX + PyY

40 = 8X + 2Y

The budget constraint is as follows:



 

b.). Slope of the budget line = "-\frac{Px}{Py} = -\frac{8}{2} = -4"

 

c.). With an increase in the price of good X, the consumer can purchase a smaller quantity of X than before. Therefore, with the rise in the price of good X, the budget line will shift to the left to the new position on the budget line.


d.). With a fall in the price of good Y, the consumer will be able to purchase more of good Y with the given budget, and therefore, the budget line will shift above to a new position.